A man rode a bicycle for 12 miles and then hiked on an additional 8 miles. The total time for the trip was 5 hr, if his rate when he was riding a bicycle was 10 miles per hour faster than his rate walking. What was each rate?: A man rode a bicycle for 12 miles and then hiked on an additional 8 miles. The total time for the trip was 5 hr, if his rate when he was riding a bicycle was 10 miles per hour faster than his rate walking. What was each rate?
Answer by checkley77(3394)

5=12/(x+10)+8/x
5=[12x+8(x+10)]/x(x+10)
5=[12x+8x+80]/x^2+10x
5=(20x+80)/x^2+10x
5x^2+50x=20x+80
5x^2+50x-20x-80=0
5x^2+30x-80=0
5(x^2+6x-16)=0
5(x-2)(x+8)=0
x-2=0
x=2 mph for the hiking rate.
10+2=12 mph for the biking rate.
proof:
5=12/12+8/2
5=1+4
5=5